Owl Shift 10/16/99

KTeV Dart (da_ktev@ktev3)
Sat, 16 Oct 1999 07:01:28 -0500 (CDT)

Shift Summary for Owl Shift 10/16/99

On Shift: TA, RP

Hours of beam available at all intensities: 7.0 hr
Hours of beam available at nominal intensity: 7.0 hr
Hours of E832 data (special runs not included): 6.5hr
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

Start new run because of L3 accept fraction error.
Changing tapes.

========================================================================
Runs ended this shift:
Run 14848 - 1.8M events - BEAM_799
Run 14850

Tapes brought to FCC this shift:
(including tapes for online split)

UPA082,83,84
UPB092,83,84
UPC082,83,84 (upc084 missing external case label - one added)
UPS028

========================================================================
Accesses during shift:

========================================================================
Overview of shift:
E.g., calibrate CsI, E832 datataking

E799 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14847 00:15 Ktevana error. kfrend.
00:45 L3 accept fraction error. Page Sasha. Fresh Start.
Problems with pipeline crate 8. Hard reboot. Still bit
problem. TCL error on next try. Page Sasha.
pipe_cleanup and Fresh Start
14849 01:30 Start Run.
02:30 DAQ_DEAD alarms. Sem ~9.4E12 (request is 9.0E12).
Call MCR to turn intensity down to 9.0E12
03:30 Ed called. At next start run he wants to install
new CsI constants since E/p is drifting to ~0.984.
05:10 Tapes Full. Paged Ed. Loaded new CsI constants.
14850 05:25 Run Started. Mounted onsplit tape upM018.
05:35 Spill 9 - updated the pmt gain reference for next run.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift